OpenStudy (anonymous):

Given that f(x)=2x+5 and g(x)=x-7, solve for f(g(x)) when x=-3

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

wait this is a little complicated

OpenStudy (anonymous):

g(1) = 2(1) - 3 = -1 So f(g(1)) = f(-1) = (-1)² + 2 = 1 + 2 = 3 f(g(1)) = 3 <==ANSWER

OpenStudy (larseighner):

Start from the inside. What is g(x) when x = -3?

OpenStudy (anonymous):

wait gracieee whats your last name?

OpenStudy (larseighner):

What problem were you trying to solve @BigDSpeedGod ?

OpenStudy (anonymous):

nothin i helped

OpenStudy (larseighner):

Your work had nothing to do with the problem stated in the question.

OpenStudy (anonymous):

yes it did

OpenStudy (anonymous):

BIG D out

OpenStudy (larseighner):

Okay the actual answer is found this way: Substitute the value -3 for x in g(x). Use that result to substitute in x in f(x). It seems to me that gives the answer as -15.